OpenStudy (satsuki):

Solve the trigonometric equation on the interval [0,2pi) 2sin^2(theta)-cos(theta)-1=0 I've got no clue how to start this, please help.

OpenStudy (anonymous):

rewrite \(\sin^2(\theta)\) as \(1-\cos^2(\theta)\) and then you have a quadratic equation with only cosines

OpenStudy (anonymous):

\[2(1-\cos^2(\theta))-\cos(\theta)-1=0\] etc you good from there?

OpenStudy (satsuki):

I should get cos^2θ-cosθ-(1/2)=0 right?

OpenStudy (satsuki):

Please help

OpenStudy (anonymous):

oh sorry lost you

OpenStudy (anonymous):

\[2(1-\cos^2(\theta))-\cos(\theta)-1=0\] \[2-2\cos^2(x)-\cos(x)-1=0\] \[2\cos^2(x)+\cos(x)-1=0\] \[(2\cos(x)-1)(\cos(x)+1)=0\]

OpenStudy (anonymous):

easier to leave the 2 there, it factors nicely

OpenStudy (anonymous):

you get \[\cos(x)=-1\] or \[\cos(x)=\frac{1}{2}\] and you can solve those

OpenStudy (satsuki):

You are amazing, thank you sincerely.